### v1

# LINKAGE: Listwise Ranking among Varied-Quality References for Non-Factoid QA Evaluation via LLMs

Abstract
Non-Factoid (NF) Question Answering (QA) is challenging to evaluate due to diverse potential answers and no objective criterion. The commonly used automatic evaluation metrics like ROUGE or BERTScore cannot accurately measure semantic similarities or answers from different perspectives. Recently, Large Language Models (LLMs) have been resorted to for NFQA evaluation due to their compelling performance on various NLP tasks. Common approaches include pointwise scoring of each candidate answer and pairwise comparisons between answers. Inspired by the evolution from pointwise to pairwise to listwise in learning-to-rank methods, we propose a novel listwise NFQA evaluation approach, that utilizes LLMs to rank candidate answers in a list of reference answers sorted by descending quality. Moreover, for NF questions that do not have multi-grade or any golden answers, we leverage LLMs to generate the reference answer list of various quality to facilitate the listwise evaluation. Extensive experimental results on three NFQA datasets, i.e., ANTIQUE, the TREC-DL-NF, and WebGLM show that our method has significantly higher correlations with human annotations compared to automatic scores and common pointwise and pairwise approaches.
LINKAGE: Listwise Ranking among Varied-Quality References for Non-Factoid QA Evaluation via LLMs
Sihui Yang1,2 Keping Bi1,2 Wanqing Cui1,2 Jiafeng Guo1,2 Xueqi Cheng1,2 1CAS Key Laboratory of Network Data Science and Technology, Institute of Computing Technology, Chinese Academy of Sciences; 2School of Computer Science and Technology, University of Chinese Academy of Sciences; {yangsihui22s, cuiwanqing18z, bikeping, guojiafeng, cxq}@ict.ac.cn,
1 Introduction
In recent years, studies on various aspects of Large Language Models (LLMs) have been drawing significant attention, a majority of which are based on the task of factoid question answering (QA) (Saad-Falcon et al., 2024; Xu et al., 2024; Lee et al., 2022). New evaluation metrics and benchmarks have also been proposed for assessing the factuality of LLMs Wang et al. (2023); Min et al. (2023). However, much less research has been conducted on non-factoid question answering (NFQA), which usually requires long-form answers to answer open-ended non-factoid questions (NFQ), such as explanations, opinions, or descriptions. This can be attributed to the inherent difficulty of the NFQA task and the lack of a well-recognized metric to evaluate the generated long-form answers. Effective evaluation of NFQA is the foundation of developing advanced techniques to enhance the quality of LLMs-generated non-factoid answers.
Evaluating NFQA is challenging since non-factoid questions often involve subjective interpretations and the potential answers can be diverse instead of a definite fact. Most prior work used automatic evaluation metrics such as measuring word overlaps (e.g., ROUGE Lin (2004) and BLEUPapineni et al. (2002)) and semantic similarities (e.g., BERTScore Zhang et al. (2019)) with the ground truth answers. To ensure the evaluation reliability, a small amount of manual annotations are also incorporated to compare the NFQA performance. However, both of them have some limitations: Automatic metrics like ROUGE, BLEU, and BERTScore cannot accurately measure the responses with semantically similar expressions or from a different but reasonable perspective respectively; Human evaluations, although more accurate in measuring various aspects of the long-form answers, often require annotators to have related knowledge to be reliable and are too expensive to apply on a large scale.Krishna et al. (2021); Liu et al. (2023). Moreover, even for humans, evaluation of NFQA can still be challenging due to the requirement of domain knowledge as well as subjective interpretations of the questions and judgment criterions.
By ingesting large-scale data from multi-tasks, LLMs, such as the GPT series, have achieved compelling performance on numerous Natural Language Processing (NLP) tasks, and sometimes even outperform humans Zhao et al. (2023). Increasing attention has been drawn to leveraging LLMs as surrogates for large-scale evaluation on model-generated responses Min et al. (2023); Saad-Falcon et al. (2024); Fu et al. (2023). Following the routines of human evaluation, approaches that leverage LLMs as judges often adopt the ways of pointwise scoring that grades each candidate answer individually and pairwise comparisons that compare pairs of answersZheng et al. (2024). The pair for comparison can be two candidate answers or a candidate answer and a ground truth answer. Figure 1 shows a concrete example of these two approaches.
Pointwise grading is hard since the accurate perception of differences between each grade can be difficult. Subtle differences between candidates may not be discerned and reflected in the final score. Pairwise comparison is relatively easier and can be more accurate but it is not scalable to the large number of candidates when the comparisons are between candidates. In contrast, there is no such issue when comparing the pair of a candidate and a ground truth answer. However, it is not feasible when the ground truth is unavailable. Moreover, when only a single ground truth exists, the evaluation may not be accurate to cover various aspects.
Inspired by the evolution of learning to rank in information retrieval, i.e., from pointwise to pairwise to listwise Liu et al. (2009); Cao et al. (2007), we propose a listwise NFQA evaluation approach that leverages LLMs to conduct ListwIse raNKing AmonG varied-quality referencEs, abbreviated as LINKAGE. Specifically, we use LLMs to assess a candidate answer by its rank in a list of reference answers sorted by quality descendingly. When there are ground truth answers of multiple grades, they can be used as the varied-quality references. When there is only one or no golden answer, we will construct some examples of multi-grade answers and utilize the in-context learning ability of LLMs to generate more reference answers of different quality. Compared to the pointwise and pairwise approach, listwise ranking can yield more accurate assessment since the LLM judge can take reference answers of various quality into consideration simultaneously. When only one reference answer is used, our method degenerates to pairwise comparisons with a ground truth answer. Additionally, given an ordered reference answer list, LLMs only ingest the reference list and candidate answer once, which costs much less than comparing each reference answer with the candidate pairwise and aggregate the score.
We conduct extensive experiments on three NFQA datasets: ANTIQUE Hashemi et al. (2020), the non-factoid portion of TREC DL Craswell et al. (2020, 2021), and WebGLM Liu et al. (2023). ANTIQUE and TREC DL have multi-grade manual annotations on the candidate answers while WebGLM is a non-factoid QA dataset based on Retrieval Augmented Generation (RAG) that provides retrieval passages and a single ground truth answer. Under the settings where there are multiple, single, or none ground truth answers, our method outperforms the automatic similarity scores, as well as pointwise, and pairwise LLM evaluation methods significantly in terms of the correlation with human judgments. By offering more accurate NFQA evaluation, our work can pave the way for future studies on improving NFQA performance, especially promoting LLMs to become more capable of answering complex questions.
2 Related Work
2.1 Non-factoid Question Answering(QA)
Non-factoid question answering (NFQA) is a complex challenge, characterized by open-ended queries that require complex responses such as descriptions, opinions, or explanations.(Yulianti et al., 2017; Cohen and Croft, 2016). These responses are usually extensive, often requiring paragraph-level answers. The most used benchmark in NFQA is the ELI5 dataset (Fan et al., 2019), which contains 272,000 questions from the "Explain Like I’m Five" Reddit forum. Moreover, multi-document NFQA datasets like WebGLM (Liu et al., 2023), WikihowQA (Bolotova-Baranova et al., 2023) integrate multiple detailed passage-level answers to form long-form answers to NFQ. ANTIQUE Hashemi et al. (2020) provides a reliable collection with complete relevance annotations of NFQA.
2.2 Non-factoid QA Evaluation
Prior NFQA approaches can be categorized into three categories:
Automatic Evaluation: Before the emergence of LLM, the most commonly used evaluation methods were automatic metrics, such as ROUGE (Lin, 2004), BLEU (Papineni et al., 2002), and BERTScore (Zhang et al., 2019). These metrics evaluate the quality of a generated answer based on text similarity between the answer and human-written answers. However, these automatic metrics calculate scores through n-gram similarity, ignoring semantic information. For instance, Krishna et al. (2021) show that ROUGE is an ineffective metric in long-form question answer tasks. Another way to implement automatic evaluation is by training a model with human evaluation preferences to conduct automatic assessment, such as QAFactEval(Fabbri et al., 2021) and RankGen(Krishna et al., 2022). However, these methods struggle to generalize to out-of-domain QA evaluation due to limited human annotations.
Human Evaluation: In NFQA tasks, human annotations are usually considered the golden standard. Hurdles (Krishna et al., 2021), WebGPT (Nakano et al., 2021), WikihowQA (Bolotova-Baranova et al., 2023) both ask human annotators to choose their preferred answer between the answer generated by the model and the golden answer. Moreover, to compensate for human lack of understanding in certain domains, they can refer to evidence documents during evaluation. However, human evaluation is expensive and therefore difficult to adopt on a large scale.
LLM Evaluation: As LLMs advance, they are gradually replacing costly human annotations. GPTScore (Fu et al., 2023) uses the generation probability of LLMs to evaluate the model-generated output. LLM-EvalLin and Chen (2023) uses a unique prompt-based evaluation method for open-domain conversations with LLMs. PRD (Li et al., 2023) and CHATEVAL (Chan et al., 2023) integrate different LLMs’ evaluation results by ranking, discussing, and debating among LLMs. The advantage of using LLMs as evaluators lies in their explainability and scalability. However, they also encounter issues such as position bias, verbosity bias, and self-enhancement bias. (Zheng et al., 2024) There is a lack of research specifically focused on LLM evaluation for NFQA.
Figure 1: Pointwise scoring evaluation, pairwise comparison evaluation and our LINKAGE evaluation approaches.
3 Method
In this section, we propose a ListwIse raNKing AmonG varied-quality referencEs method (LINKAGE) for evaluating NFQA. We formally define the task of NFQA evaluation and introduce some basic evaluation approaches, then introduce the details of our LINKAGE.
3.1 Preliminary
Task Definition: Given a non-factoid question $q$ and its corresponding n candidate answers $\mathcal{C}=\{c_{1},c_{2},...,c_{n}\}$ to be evaluated, where $c_{i}$ represents the $i$-th candidate answer. The goal is to score each answer with a scorer $Score(c_{i})$. The ground truth set of $q$ is $\mathcal{G}=\{g_{1},g_{2},...,g_{k}\}$, in which $g_{i}$ represents the $i$-th ground truth. In this paper, the scorer is LLM and we use a prompt $\mathcal{P}$ to query the LLM to get the scoring results.
Currently, the commonly used scoring methods based on LLM are pointwise and pairwise approachesZheng et al. (2024).
Pointwise Evaluation: The pointwise evaluation approach assesses an answer $c_{i}$ only based on its relevance and quality regarding the question $q$. As shown in Figure 1, the evaluation process may be conducted with or without using ground truth answers as references.
$Score_{\mathrm{point}}(c_{i})=f(\mathcal{P}_{\mathrm{point}},q,c_{i},\mathcal{R}),$ (1)
in which $f(P_{\mathrm{point}},\cdot)$ represents querying the LLM through prompt $\mathcal{P}_{\mathrm{point}}$. $\mathcal{R}=[r_{1},r_{2},\ldots,r_{m}]$ is a reference answer list sorted by quality in descending order, which can be $\mathcal{G}$, a subset of $\mathcal{G}$, or $\emptyset$.
Pointwise grading is easy to conduct but difficult to accurately perceive grade differences. The subtle differences among candidates may not be distinguished and reflected in the final score.
Pairwise Evaluation: As shown in Figure 1, the pairwise evaluation approach performs a pairwise comparison between answers. The pairs can be two candidate answers,
$Score_{\mathrm{pair}}(c_{i})=\sum_{c_{j}\in\mathcal{C}\setminus\{c_{i}\}}f(\mathcal{P}_{\mathrm{pair}},q,c_{i},c_{j}).$ (2)
However, the number of comparisons between candidate answer pairs grows exponentially with the number of candidate answers, and thus cannot be scaled to a large number of candidates. The pair can also be a candidate answer and a reference answer,
$Score_{\mathrm{pair}}(c_{i})=\sum_{r_{j}\in\mathcal{R}}w_{l_{j}}*f(\mathcal{P}_{\mathrm{pair}},q,c_{i},r_{j}),$ (3)
$f(\mathcal{P}_{\mathrm{pair}},q,c_{i},r_{j})=\left\{\begin{aligned} 1&\text{, if }c_{i}\text{ is better}\\ -1&\text{, if }r_{j}\text{ is better}\\ 0&\text{, otherwise}\\ \end{aligned}\right..$ (4)
$\mathcal{R}$ can be $\mathcal{G}$ or a subset of $\mathcal{G}$. $w_{l_{j}}$ is the weight corresponding to certain grade $l_{j}$ of answer $r_{j}$. In this way, the pairwise approach scores a candidate answer by comparing it with each answer in the reference answer list.
Pairwise comparison is relatively easier and can be more accurate, but when there is only a single ground truth, evaluation becomes less accurate because it is difficult for a single ground truth to cover various aspects of NFQA
3.2 Listwise Ranking Evaluation (LINKAGE)
Figure 1 shows how our LINKAGE works. Specifically, given a reference answer list sorted by descending quality and the answer to be evaluated, the scorer judges its quality by deciding where it should be ranked among the reference answer list,
$Score_{\mathrm{pair}}(c_{i})=f(\mathcal{P}_{\mathrm{list}},q,c_{i},\mathcal{R}).$ (5)
The higher the ranking, the better the quality.
Please note the difference between our method and the pointwise approach with references. Although both methods ask LLMs to directly output a numerical value, in the pointwise approach, references are used to provide a criterion for scoring, and the assignment only focuses on the quality of $c_{i}$ itself rather than comparisons. The listwise ranking approach relies on comparing it with all reference answers to determine where the answer should be ranked.
3.3 Reference List Construction
Reference answer list $\mathcal{R}$ in LINKAGE is composed of multiple answers ordered in descending quality. Compared to providing LLMs with only one ground truth, more references with different styles and quality enable the LLM evaluators to learn implicit evaluation guidelines from $\mathcal{R}$. The collection method of $\mathcal{R}$ depends on the composition of the ground truth set of the dataset, and we discuss it in three situations:
3.3.1 Multi-grade Ground Truth
When multiple grades of ground truth answers are available, references can be sampled directly from these answers. For instance, ANTIQUE and TREC DL contain multiple answers annotated with four relevant labels.
To reduce bias and ensure the reliability of evaluation results, we randomize the sampling process multiple times. Additionally, the length and the distribution of $\mathcal{R}$ also impact the results. We discuss this in detail in Section 5.2.
3.3.2 Single-grade Ground Truth
Some NFQA datasets, such as WebGLM, only contain a single grade of ground truth. For this scenario, we prompt LLMs to generate answers of varying quality to serve as references. Specifically, we first prompt LLMs to answer the question based on the original golden answer, thus obtaining a new high-quality golden answer. The prompt is in Figure 8 (Appendix A.2). This step ensures that both the golden reference and other reference answers are generated by LLMs, avoiding the introduction of style bias between human and machine writing. We then use the prompt in Figure 9 (Appendix A.2) to obtain other lower-quality reference answers. To ensure the diversity of references, we use three LLMs to generate separate lists of reference answers. Then we randomly sample reference answers from three lists to form $\mathcal{R}$ for each grade.
3.3.3 Absence of Ground Truth
In real-world scenarios, non-factoid questions may not have reference answers. To tackle the problem of ground truth missing, considering the powerful capabilities of LLMs like GPT-4 OpenAI (2022a), we get a quality-assured answer from GPT-4 directly. The ways of generating reference answers of other quality are the same as described in Section 3.3.2.
4 Experimental Settings
4.1 Datasets
We evaluate the effectiveness of baseline methods and our proposed LINKAGE using the following three datasets.
ANTIQUE Hashemi et al. (2020) dataset contains 2,626 open-domain non-factoid questions asked by real users in a community question answering service, i.e., Yahoo! Answers. Similar to TREC-DL, all passages are graded into four levels (3: reasonable and convincing, 2: not sufficiently convincing, 1: unreasonable, 0: make no sense). We merge the 200 questions from the test set and the 300 questions randomly sampled from the training set, yielding a total of 500 queries as our experiment dataset.
Table 1: Statistics of ANTIQUE and TREC-DL-NF we use in experiments.
Number statistics ANTIQUE TREC-DL-NF
#Question 500 55
#Avg doc labeled 3 5.8 9.6
#Avg doc labeled 2 4.5 18.1
#Avg doc labeled 1 6.5 24.9
#Avg doc labeled 0 3.6 48.0
#Avg total documents 20.4 100.7
TREC-DL-NF Craswell et al. (2020, 2021) In our experiments, we use TREC-DL 2019, 2020 datasets, which comprise 43 and 54 MS MARCO queries respectively. Each question has multiple passages labeled with four levels of relevance (3: perfectly relevant, 2: highly relevant, 1: related, 0: irrelevant). Not all questions are NF questions, so we filter factoid questions with a non-factoid question category classifier Bolotova et al. (2022). This leaves us a total of 55 non-factoid questions, denoted as TREC-DL-NF.
The statistics of ANQIQUE and TREC-DL-NF can be found in Table 1.
WebGLM Liu et al. (2023) is a high-quality quoted long-formed retrieval-augmented QA dataset. Each question is accompanied by 5 top-ranked documents retrieved by a vanilla Contriever Izacard et al. (2021). Question and corresponding candidate references are fed together to OpenAI text-davinci-003 Ye et al. (2023) to generate long-formed answers by 1-shot in-context learning. To obtain candidate answers of different styles and quality, we use gpt-3.5-turbo-16k  OpenAI (2022b) to generate two answers with 5 relevant and 3 relevant plus 2 irrelevant documents respectively. The third answer is generated by Mistral-7B-Instruct-v0.2 Jiang et al. (2023) with 5 relevant documents. We sample 50 cases and manually label three candidate answers with three levels (3,2,1). Details about manual annotation are in the Appendix D.
4.2 Methods for Comparison
We compare the following NFQA evaluation baselines and our LINKAGE under different situations.
4.2.1 Baselines
Automatic Metrics:
ROUGELin (2004), BERTScoreZhang et al. (2019), BLEUPapineni et al. (2002) are all reference-based metrics based on text similarity. ROUGE and BLEU focus on exact n-gram matching, while BERTScore evaluates the semantic similarity of embeddings.
LLM Evaluation Baselines:
  • PointwiseR=∅: This method asks LLMs to directly assign a quality score from 1 to 10 to the candidate answer without any reference answers.
  • PointwiseR≠∅: Based on the basic pointwise method, this method also provides a list of reference answers sorted in descending order of quality for LLMs to refer to when scoring.
  • Pairwise: This method scores a candidate answer based on comparing it with each answer in the reference list. To eliminate position bias, i.e., the LLM judge might favor the forward-positioned one when comparing two answers, we randomly permute the positions of the candidate answer and ground truth answer during evaluation.
Table 2: The performance of different methods on ANTIQUE and TREC-DL-NF. K, S, and P represent Kendall’s tau, Pearson’s r, and Spearman’s rho coefficient respectively. The best results of each evaluator model are in bold.
Method ANTIQUE TREC-DL-NF
K S P K S P
Automatic Metrics ROUGE-1 0.2088 0.2563 0.2878 0.2442 0.3060 0.3412
ROUGE-2 0.1807 0.2089 0.2281 0.2064 0.2441 0.2808
ROUGE-L 0.2012 0.2463 0.2708 0.2171 0.2721 0.3178
BERTScore 0.1562 0.1938 0.1950 0.2258 0.2824 0.2842
BLEU 0.1808 0.2153 0.2063 0.2106 0.2650 0.2208
LLM Evaluation on Mistral PointwiseR=∅ 0.2202 0.2499 0.2519 0.2366 0.2773 0.2677
PointwiseR≠∅ 0.2229 0.2516 0.2547 0.3138 0.3382 0.3302
Pairwise 0.1827 0.2134 0.2132 0.2501 0.2967 0.2939
LINKAGE on Mistral LINKAGE0_shot 0.3585 0.3790 0.3893 0.3287 0.3539 0.3401
LINKAGEfew_shot 0.3742 0.4200 0.4373 0.4312 0.4725 0.4958
LLM Evaluation on ChatGPT PointwiseR=∅ 0.2777 0.3118 0.3244 0.3176 0.3640 0.3660
PointwiseR≠∅ 0.2752 0.3112 0.3224 0.3746 0.4288 0.4449
Pairwise 0.2979 0.3494 0.3756 0.3204 0.3692 0.3749
LINKAGE on ChatGPT LINKAGE0_shot 0.3070 0.3404 0.3514 0.3923 0.4315 0.4376
LINKAGEfew_shot 0.3096 0.3543 0.3688 0.3993 0.4325 0.4481
Table 3: Results for the situation of single-grade ground truth. The best results of each model are in bold.
Model Method ANTIQUE TREC-DL-NF
K S K S
[…]
Pairwise1GT 20.89 23.41 30.43 36.62
[…]
Pairwise1GT 29.88 32.32 30.28 34.14
[…]
Table 4: Results for the situation of absence of ground truth. The best results of each model are in bold.
Model Method ANTIQUE TREC-DL
K S K S
[…]
4.2.2 LINKAGE
LINKAGE: To ensure that $\mathcal{R}$ uniformly contains answers of varying quality, we randomly select the same number of reference answers from the answer set of each level to create the reference answer list. For TREC-DL-NF, the grades of answers in the reference list are $\mathcal{L}=(3,2,1,0)$. For ANTIQUE, $\mathcal{L}=(3,3,2,2,1,1,0,0)$, which are intuitively reasonable and balanced settings.
LINKAGE-1GT: We also test the case where there is only one ground truth. For questions with multi-grade answers, we randomly sample one answer from the highest-grade ground truth set as the only ground truth to simulate this situation.
LINKAGE-0GT: In this case, we do not use any labeled ground truth to simulate the situation where no ground truth is available.
4.3 Evaluation Metrics
To evaluate the effectiveness of NFQA evaluation, we use Kendall’s tau, Pearson’s r and Spearman’s rho coefficient to calculate the extent of consistency between the resulting sorted sequences and the manually labeled sequences. Spearman’s rho coefficient is chosen as our primary metric due to its balance between robustness and sensitivity to monotonic relationships.
4.4 Implementation Details
The evaluation experiments are based on two representative LLMs: (i) The open-source model Mistral (Mistral-7B-Instruct-v0.2  [^1] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2Jiang et al. (2023). (ii) The close-source model ChatGPT (gpt-3.5-turbo-16k) OpenAI (2022b), for which results are obtained through API. The temperature for all experiments is set to 0.8.
When only one or no ground truth exists, we use gpt-4-1106-preview OpenAI (2022a) to generate the golden answer. For generating other references with descending quality, we use three different LLMs in 3-shot setting: (i) Mistral-7B-Instruct-v0.2, (ii) gpt-3.5-turbo-16k, (iii) Meta-Llama-3-8B-Instruct  [^2] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct Meta (2024). All our experiments are done on a single Tesla A100 80G GPU.
5 Experimental Results
Figure 2: An example of our LINKAGE  compared with pointwise and pairwise approaches. We standardized the score range of all methods to $[0,10]$ for easy comparison and understanding.
Table 5: Results on WEBGLM based on Mistral. RL, BS, and B represent ROUGE-L, BERTScore, and BLUE, respectively. Acc(b) means the accuracy of finding the best answer. Acc(b+w) means the accuracy of finding both the best and the worst answers.
RL BS B PointR≠∅ Pair LINKAGE
Acc(b) 0.42 0.48 0.50 0.46 0.54 0.76
Acc(b+w) 0.32 0.38 0.44 0.22 0.32 0.34
Table 6: Different composition of $\mathcal{R}$ on ANTIQUE and TREC-DL-NF and using Mistral. The settings we use in LINKAGE are in bold.
Dataset $\mathcal{R}$| $\mathcal{R}$ 0-shot 3-shot
K S K S
ANTIQUE 4 3210 24.68 25.95 26.79 29.65
8 33321000 29.88 32.62 31.54 34.81
33221100 21.40 22.93 25.01 27.18
32221110 25.77 29.16 28.10 29.65
TREC-DL-NF 4 3210 25.79 27.52 31.78 34.83
8 33321000 24.03 25.81 30.90 34.24
33221100 25.85 27.61 32.70 36.30
32221110 24.94 27.01 30.97 34.39
5.1 Overall Results
The results on the multi-grade ground truth situation, single-grade ground truth situation, and absence of ground truth situation are shown in Table 2, Table 3, and Table 4 respectively. The results on WebGLM are shown in Table 5. It can be seen that our method always shows better consistency with human evaluation.
Additionally, we have the following observations:
LLM-based methods perform generally better than automatic metrics. This indicates that automatic metrics have limitations in NFQA evaluation, therefore should be used with caution in future research. Among LLM-based methods, our proposed LINKAGE outperforms all other baselines by a significantly large margin leveraging both Mistral and ChatGPT. This confirms the superiority of listwise approach over the pointwise and pairwise approaches on NFQA evaluation.
Few-shot in-context Learning can enhance the performance of LINKAGE. Comparing with results under few-shot and zero-shot, providing LLMs with a few examples can help demonstrate the evaluation task more clearly. Compared to Mistral, the enhancement of few-shot ICL on ChatGPT is less. We think that it is because ChatGPT has a much better understanding of instructions so the few-shot example does not help it much. More experiment details about few-shot learning are in the Appendix B.1.
Reference answer list is important for understanding NFQ evaluation criteria. By analyzing the pointwise method results with and without reference, we find PointwiseR≠∅ always performs better. In some cases, it can even exceed the performance of pairwise methods. This indicates that providing the reference answer list helps LLMs understand NFQ evaluation criteria so that PointwiseR≠∅ can assign a more reliable score than PointwiseR=∅. This further illustrates that providing $\mathcal{R}$ in evaluating NFQA can lead to significant gains.
LINKAGE is applicable in various of situations. Table 3 and Table 4 show that LINKAGE-1GT and LINKAGE-0GT both perform the best among all LLM evaluation methods. This illustrates that our method is still effective when generalized to other evaluation scenarios, i.e., when there is only one ground truth or no ground truth.
5.2 Study on the Reference List Composition
We conduct experiments on different reference distributions to analyze their impact. As shown in Table 6, varying length and distribution of $\mathcal{R}$ affects the performance of LINKAGE. To ensure the fairness of the experiment, candidate answers are the same for each setting.
The impact of length depends on the quality of the dataset. ANTIQUE is collected from web data and contains more noise, so increasing the number of references can help LLMs better build evaluation criteria. The conclusion on TREC-DL-NF is the opposite. For quality assurance datasets, increasing the number of references, however, exacerbates the burden of understanding long texts, thereby impairing evaluation performance. For the grade distribution of reference answers, uniform sampling always brings the best results, as it allows LLMs to understand all grades of answers while avoiding introducing grade preference bias.
5.3 Study on the Reference List Randomness
Our experiments involve random sampling of the ground truth set, so we evaluate the results under 3 randomizations to analyze the impact of randomness on performance. Details about experiments are in Appendix B.2. From Figure 3, we can observe that for all LLMs on all datasets, the standard deviations of the experiments are always small. This indicates that the randomness of the selection of reference answers has little impact on the evaluation results, which proves that the improvement brought by our method is significant.
Figure 3: Comparison of Spearman Correlation for Mistral and ChatGPT on ANTIQUE and TREC-DL-NF. The error bars denote the standard deviation, illustrating the variability in the results.
6 Case Study
We conduct case studies to qualitatively compare the results of different methods. As shown in the Figure 2, because candidate answer 1 contains many matching keywords, even though it does not effectively answer the question, pointwise method and pairwise method both assign it a high score. As a result, the two candidate answers cannot be effectively distinguished. In contrast, our LINKAGE can better distinguish the fine-grained quality differences between candidate answers and obtain results that are more consistent with humans.
7 Conclusion
In this paper, we propose a listwise NFQA evaluation approach (LINKAGE), which leverages LLMs to assess a candidate answer by its rank in a list of sorted reference answers. Our approach is capable of considering reference answers of various quality simultaneously. Therefore, it can enable LLMs to establish a better evaluation system and yield more accurate assessments. Extensive experiments on three datasets, i.e., ANTIQUE, TREC-DL-NF, and WebGLM, demonstrate the effectiveness of our method, whether it is in situations with multi-grade ground truth answers, single-grade ground truth answers, or no ground truth. Hoping this more accurate evaluation method can promote future research on NFQA.
Limitations
There are two primary limitations: (i) Our method demands multiple grading labels when constructing the reference answer list. When grading labels are missing, utilizing LLMs to generate reference answers increases the cost of inference. How to reduce the computational cost requires further research in the future. (ii) Compared with the pointwise and pairwise methods, the listwise method considers the relationship between all documents, so it requires the scoring model to have a good long-text understanding ability.
Appendix A Instruction Details
A.1 Instruction for Evaluation
The evaluation prompts are adopted in LINKAGE and LLM baselines (detailedly introduced in Sec4.2). These prompts are fed to LLMs, allowing them to generate scores, preferences or rankings.
Figure 4: Instruction for pointwise scoring without references.
Figure 5: Instruction for pointwise scoring with references.
Figure 6: Instruction for pairwise comparison.
Figure 7: Instruction for our proposed LINKAGE.
A.2 Instruction for Generating Reference List
Figure 8: Instruction for generating the highest standard reference answer.
Figure 9: Instruction for generating other reference answers in $\mathcal{R}$ sorted by quality descendingly.
Appendix B Experiment Details
B.1 N-shot LINKAGE performance
We conduct several sets of few-shot experiments of LINKAGE on TREC-DL-NF and ANTIQUE using Mistral. Results are in Table 7 and Table 8. We find the number of samples cannot be too large. When the number exceeds a certain value, the performance will deteriorate. This is because the shot number increasing leads to a significant increase in the input length, which will make the LLMs difficult to understand.
n-shot Kendal Spearman Pearson
$n$=0 0.3287 0.3539 0.3401
$n$=1 0.4246 0.4656 0.4724
$n$=3 0.4312 0.4752 0.4958
$n$=5 0.4339 0.4725 0.4958
Table 7: The performance of LINKAGE under different few shot setting on TREC-DL-NF using Mistral.
n-shot Kendal Spearman Pearson
$n$=0 0.2900 0.3101 0.3172
$n$=1 0.3850 0.4183 0.4256
$n$=3 0.3696 0.4012 0.4122
$n$=5 0.3654 0.3934 0.4041
Table 8: The performance of LINKAGE under different few shot setting on ANTIQUE using Mistral.
B.2 Experiments on Randomness of Reference List
Three independent experiments on randomly selecting $\mathcal{R}$ on TREC-DL-NF and ANTIQUE using ChatGPT and Mistral are shown in Table 9 and Table 10.
Table 9: Average Spearman coefficient and standard deviation of randomly selecting $\mathcal{R}$ in three dependent experiments on TREC-DL-NF using Mistral and ChatGPT.
Model Method Spearman 1 Spearman 2 Spearman 3 Average Std
Mistral PointwiseR≠∅ 0.3382 0.3463 0.3567 0.3471 0.0093
Pairwise 0.2967 0.2783 0.2912 0.2887 0.0094
LINKAGEfew_shot 0.4725 0.4579 0.4520 0.4608 0.0105
ChatGPT PointwiseR≠∅ 0.2777 0.4288 0.4526 0.3864 0.0948
Pairwise 0.3692 0.3687 0.3544 0.3641 0.0083
LINKAGEfew_shot 0.4094 0.3854 0.4325 0.4091 0.0235
Table 10: Average Spearman coefficient and standard deviation of randomly select $\mathcal{R}$ in three dependent experiments on ANTIQUE using Mistral and ChatGPT.
Model Method Spearman 1 Spearman 2 Spearman 3 Average Std
Mistral PointwiseR≠∅ 0.2516 0.1781 0.1778 0.2025 0.0425
Pairwise 0.2210 0.2059 0.2134 0.2134 0.0082
LINKAGEfew_shot 0.4200 0.4078 0.4122 0.4133 0.0062
ChatGPT PointwiseR≠∅ 0.3118 0.3250 0.3180 0.3182 0.0094
Pairwise 0.3495 0.3387 0.3402 0.3428 0.0076
LINKAGEfew_shot 0.3543 0.3143 0.3339 0.3340 0.0283
Appendix C Case Study
The details of case in Section 6 of the main paper is in Figure 10.
Figure 10: An example of our LINKAGE  compared with Pointwise and Pairwise approach.
Appendix D Human Annotation
We recruit one domain expert who has earned at least a bachelor’s degree in Computer Science to annotate WEBGLM candidate answer’s quality label. The instruction is shown in Figure 11.
Figure 11: Instructions for labeling WEBGLM for human annotators.